part of a fuel moisture formula I am trying to solve.. the formula is:
MC=351/e^(DC/390)
I have made a physical observation determing the value of MC but now I need to solve for DC. Any help will be appreciated.
Dan Morrison
morrisondan@shaw.ca

Answer by Alan3354(69443)

part of a fuel moisture formula I am trying to solve.. the formula is:
MC=351/e^(DC/390)
I have made a physical observation determing the value of MC but now I need to solve for DC.
--------------------
MC=351/e^(DC/390)
e^(DC/390) = 351/MC
DC/390 = ln(351/MC)
DC = 390*ln(351/MC)